[figure: 9.7 cm]

MAGNETIZATION and X-RAY PHOTOEMISSION SPECTRA OF

UFe_{5}Sn SINGLE CRYSTAL

A. Bajorek^{a}, G. Chełkowska^{a}, R. Troć^{b}

aA. Chełkowski Institute of Physics, Silesian University, Uniwersytecka 4, 40-007

Katowice, Poland

bW. Trzebiatowski Institute of Low Temperature and Structure Research, Polish Academy of Sciences, P.O. Box 1410, 50-950 Wrocaw, Poland

Recently a new compound of the composition UF e_{5}Sn has been synthesised and its crystal structure recognized [1]. It crystallizes in the orthorhombic CeCu_{5}Au− type structure. UF e_{5}Sn is a ferromagnet with T_{C} = 248K [1]. Here we present for this compound the magnetic properties determined on a single crystal and the electronic structure studied by X-ray photoemission spectroscopy (XPS). A complex magnetization behaviour has been found depending on the crystal orientation. It appears that the VB spectrum of the XPS can be mainly determined by the Fe(3d) and U(5f) bands. The position of the maximum in the VB spectrum of UF e_{5}Sn is shifted by 1.0 eV in the binding energy as compared to the VB of pure U.

[1] A.P.Goncalves et al. J.Solid State Chem. 154 (2000) 551.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : poster

Corresponding author : G. Chełkowska

Address for correspondence : Institute of Physics, University of Silesia, Uniwersytecka 4, 40-007 Katowice, Poland

Email address : gchelkow@us.edu.pl